20. Determine the truth value of each statement if the domain consists of all real numbers. (4 pts.)

a) βˆƒπ‘₯(π‘₯3 = βˆ’1) b) βˆƒπ‘₯(π‘₯4 < π‘₯2) c) βˆ€π‘₯((βˆ’π‘₯)2 = π‘₯2) d) βˆ€π‘₯(2π‘₯ > π‘₯)


1
Expert's answer
2021-10-29T00:49:02-0400

a) "\exists x" ("x^3=-1)"

This is true, "x=-1"

b) "\exists x\ (x^4<x^2)"

This is true, "x=\tfrac{1}{2}" and we have "x^4=\tfrac{1}{16}<\tfrac{1}{4}=x^2"

c) "\forall x\ ((-x)^2=x^2)"

This is true, "(-x)^2=(-x)\cdot(-x)=(-1)\cdot(-1)\cdot x^2=x^2"

d) "\forall x\ (2x>x)"

This is false, if we have "x=-1" , then "2x=-2<-1=x"